Equilateral triangle Find the resultant force on the 7uC charge Graded classwork: Find the electric field at the position of 3 HC charge generated by other 2 m10 minutes 7 uC two.

solution and explanation please!
Answer #1

Force on 7\muC due to -3\muC charge is given by 9*109*7*10-6*3*10-6/22 = 0.04725 N (towards -3\muC charge)

Force on 7\muC due to 5\muC charge is given by 9*109*7*10-6*5*10-6/22 = 0.07875 N (away from 5\muC charge)

Total horizontal force on 7\muC is 0.07875-0.04725*cos60 = 0.055125 N (towards right)

Total vertical force on 7\muC is 0.04725*sin60 = 0.0409 N (vertically upwards)

Resultant force magnitude = sqrt(0.0551252+0.04092)=0.0686 N